Help Please!
It has been estimated that with a sophisticated mathematical strategy, a player can increase the probability of winning at black jack to 10/19. If you play with this strategy, what is the probability that you will win at least 8 games if you play 20? (Use normal approximation)
Thank you

Answer by stanbon(57984)

It has been estimated that with a sophisticated mathematical strategy, a player can increase the probability of winning at black jack to 10/19. If you play with this strategy, what is the probability that you will win at least 8 games if you play 20? (Use normal approximation)
------------
mean = np = 20(10/19) = 10.526
std = sqrt(npq) = sqrt[20(10/19)(9/19)] = 2.233
----
z(7.5) = (7.5-10.526)/2.233 = -1.355
z(20.5) = (20.5-10.526)/2.233 = 4.467
=======
P(8<= x <=20) = P(-1.355<= z <= 4.467) = 0.9123
===============
Cheers,
Stan H.